OpenStudy (anonymous):

I know this seems like a stupid question but how would you change 1/50 to a percent?

OpenStudy (lncognlto):

Multiply it by 2 to get 2/100. Now it is 2%.

OpenStudy (anonymous):

yup 2%

OpenStudy (anonymous):

I got dividing it by 100 but I figured that wasn't right

OpenStudy (lncognlto):

You don't literally divide 2 by 100. A whole consists of 100 parts, so a number over a hundred = that number in percentage.

OpenStudy (anonymous):

the question says to give the answer to the nearest hundredth of a percent

OpenStudy (anonymous):

yes so it is 2% or .02

OpenStudy (lncognlto):

Well, it wouldn't be 0.02 of a percent, though.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I didnt put percent where did you get that?

OpenStudy (lncognlto):

I didn't say you did, but KJ's question asks for the answer to the nearest hundredth of a percent.

OpenStudy (anonymous):

In the United States approximately one child in every 10,000 is born with PKU a syndrome that affects individuals homozygous for the recessive allele. given this incidence what percent of the population are heterozygous carriers of the recessive PKU allele? give your answer to the nearest hundredth of a percent

OpenStudy (anonymous):

BTW you can just call me Josie

OpenStudy (anonymous):

it's 1/50 not 11

OpenStudy (lncognlto):

Sorry, yes, typo.

OpenStudy (anonymous):

ok I'm trusting you. I have way too many other problems to do. I have like 30 pages due on the 8th. only have 1 done

OpenStudy (lncognlto):

Well, 1/50 is most definitely = to 2%. The question might just be messing with your head when it says to the nearest hundredth of a percent.

OpenStudy (lncognlto):

Way one: Divide the numerator by the denominator, and multiply it by 100 In your question: 1/50 = 0.02. 0.02 * 100 = 2. Way two: Multiply the numerator and the denominator by a number that will make the denominator equal 100. In your question: 1/50: multiplying 50 by 2 would make the denominator equal 100. What ever you do to the denominator, you must do to the numerator. 1*2 / 50*2 = 2/100 When a fraction has a denominator of 100, the numerator is the percent: 2.

OpenStudy (anonymous):

but I posted this question a while back and the person that answered gave me like 1.98%

OpenStudy (lncognlto):

Well, then, I would go with that. It sounds closer to the answer the question expects :)

OpenStudy (anonymous):

but then I found this http://www2.kumc.edu/genetics/risk/hardy.html

OpenStudy (lncognlto):

Ok, having a look...

OpenStudy (anonymous):

I'm so confused. there's so many different answers and I want to get this right

OpenStudy (lncognlto):

Now, I would go with 2%, as the genetic example agrees.

OpenStudy (anonymous):

I don't really think the guy knows what he is talking about. but the website doesn't really give a percent related to the question

OpenStudy (anonymous):

this is what he gave me to solve for q, you take the squared root of both sides: q 2 =1 10000 →q=1 10000 − − − − − − √ →q=0.01 so p=1-q = 0.99 2ab is the percentage of heterozygous individuals (i.e. carriers) 2(0.99)(0.01)= 0.0198 1.98 %

OpenStudy (lncognlto):

Ok, I am just going to try to work it out quickly, one sec

OpenStudy (anonymous):

ok

OpenStudy (lncognlto):

Ok, just checking the values, p and p^2 = 1, and q^2 = 1/10000, correct? Just double checking my answer.

OpenStudy (anonymous):

I don't have anything for p^2 but everything else looks right

OpenStudy (lncognlto):

Well, now I'm just confused too. However, on slide 33/34 of this, http://www.slideshare.net/zna79/population-genetics-aqa, the percentage is worked out to be 1.98%

OpenStudy (anonymous):

I am very confused myself

OpenStudy (lncognlto):

:) - Sorry I can't clear up the confusion.